Should print DRAFT over all pages, quite dark.

\usepackage[landscape]{draftcopy}

Test 11 This is a test page with landscape option 1

Test 11 This is a test page with landscape option 2

This is a testpage 1 This is a testpage 2 This is a testpage 3 This is a testpage 4